) Find the intersection points of the graphs of the given equations :
6x – y = 2 and y = 4x + 7

Answer by stanbon(75887)

Find the intersection points of the graphs of the given equations :
6x – y = 2 and y = 4x + 7
-----
y = 6x - 2
y = 4x + 7
====
Substitute and solve for "x":
6x-2 = 4x+7
2x = 9
x = 4.5
============
Solve for "y":
y = 4x + 7
y = 4(4.5) + 7 = 25
======================
Cheers,
Stan H.
===============